OpenStudy (anonymous):

5.625 mg converted to g

OpenStudy (anonymous):

Let's go thru this together. How many milligrams makes a gram?

OpenStudy (anonymous):

1 mg = .001 g

OpenStudy (anonymous):

ok that's a correct statement but I was looking for: 1000mg = 1g

OpenStudy (anonymous):

So let's find the right conversion: \[5.65mg * \frac{ 1g }{ 1000mg }\] Note that the 'mg' in the numerator cancels the 'mg' in the denominator leaving only 5.65/1000g = ?

OpenStudy (anonymous):

i got .005625 g is that correct

OpenStudy (anonymous):

see the original question you are missing a number i believe, in your equation.

OpenStudy (anonymous):

yes u r correct. yes i missed a number. Do u understand what I did?

OpenStudy (anonymous):

sure no problem

OpenStudy (anonymous):

Do u have any other questions?

OpenStudy (anonymous):

yes i do thank u. 4.5 mcL converted to daL

OpenStudy (anonymous):

mcL - micro-liter: 1mcL = 10^-6 L daL - deca-liter: 1daL = 1L Therefore, 1mcL = 10^-7 daL Can u find it from here?

OpenStudy (anonymous):

no I am stuck on this one

OpenStudy (anonymous):

Answer: \[4.5mcL * \frac{ 1daL }{ 10^7mcL }\]

OpenStudy (anonymous):

What do u get?

OpenStudy (anonymous):

1 divided by 10,000,000 x 4.5 is that correct

OpenStudy (anonymous):

i got .00000045

OpenStudy (anonymous):

Yes that is correct good job. Don't forget ur unit.

OpenStudy (anonymous):

yes. daL thank u

OpenStudy (anonymous):

Sure u r welcome

OpenStudy (anonymous):

can u help with 1 more

OpenStudy (anonymous):

5.321 ml convert to Kl

OpenStudy (anonymous):

1000mL = 1L Therefore: 1000000mL = 1KL Try to do the conversion from the info above.

OpenStudy (anonymous):

.000005321 Kl

OpenStudy (anonymous):

Good job.

OpenStudy (anonymous):

thank you so much happy thanksgiving

OpenStudy (anonymous):

No problem. Thanks and the same to u :)